Circular No. 8544

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1663 AQUILAE = POSSIBLE NOVA IN AQUILA M. Dennefeld and F. Ricquebourg, Institut d’Astrophysique de Paris;

and Y. Damerdji, Observatoire de Haute-Provence (OHP), write that they obtained a calibrated spectrum (range 465–830 nm, resolution 0.6 nm) on June 11.1 UT of the possible nova reported on IAUC 8540, using the 1.93-m telescope (+ Carelec spectrograph) at OHP. The spectrum shows a very red continuum but no sign of molecular absorption bands, so reddening appears to be present; there are, however, various metallic stellar absorption features present, compatible with an early K-type star, so this may be a symbiotic nova (though the term ‘peculiar nova’ is perhaps more appropriate). There are also strong emission lines with narrow P-Cyg profiles at Hα and O I (777 nm); the expansion velocities from these lines are 700 km/s in both cases (suggestive of a slow nova). Faint Hβ and numerous faint Fe II lines, all with P-Cyg profiles, are also seen in the range 480–640 nm. Due to the peculiar nature of this nova, continuous photometric and spectroscopic observations are encouraged.

N. N. Samus and E. Kazarovets, Institute of Astronomy, Russian Academy of Sciences, inform us that the designation V1663 Aql has been given to this nova (cf. IAUC 8540).

SUPERNOVA 2005cl IN MCG –01-53-20 Further to IAUC 8542, H. Pugh and W. Li report the LOSS discovery

of an apparent supernova on unfiltered KAIT images taken on June 12.48 (at mag 18.3) and 14.45 UT (mag 18.2). An image taken on June 2.47 also showed a marginal detection of the new object (limiting mag ∼ 18.8). SN 2005cl is located at α = 21^{h}02^{m}02^{s}.35, δ = −6^{o}17′35′′.7 (equinox 2000.0), which is 18′′.0 west and 13′′.3 north of the nucleus of MCG –01-53-20. KAIT images taken on May 12.50 (limiting mag ∼ 19.5) and 23.49 (limiting mag ∼ 19.0) showed nothing at this position.

M. Modjaz, R. Kirshner, and P. Challis, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 350–740 nm) of SN 2005cl, obtained on June 13.43 UT by P. Berlind with the F. L. Whipple Observatory 1.5-m telescope (+ FAST), shows it to be a type-IIn supernova. The spectrum consists of a blue continuum, indicating young age, and Balmer emission lines with a narrow Hα component of ∼ 500 km/s (FWHM) superposed on a weaker and broader component of ∼ 4000 km/s (FWHM).

2005 June 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green